# Wave-like amplification of near-threshold two-particle reactions: from muon-catalyzed fusion to $\Lambda\bar{\Lambda}$ production at $e^{-}e^{+}$ annihilation

### v2

@@ Abstract
The effect of wave-like enhancement of the quantity $\sigma_{0}(E)v$ as $E\rightarrow 0$ was found by us in [1] in analysing fusion reactions (6) and (7). In the present work, we have investigated its manifestation in the experimental cross sections of $\Lambda\bar{\Lambda}$ production at $e^{-}e^{+}$ annihilation; however, in our opinion, it should be taken into account in any two-particle near-threshold reaction. It seems to us that a natural further development of the approach would be its generalization to describe the reaction of $\Lambda_{c}\bar{\Lambda}_{c}$ pair production at $e^{-}e^{+}$ annihilation [20]. However, it is necessary to include the Coulomb interaction between $\Lambda_{c}$ and $\bar{\Lambda}_{c}$ in the computational scheme , as well as to take into account the d-wave (3,4). The obtained formulas could be useful for analyzing the oscillatory behavior of the electromagnetic form factors of hyperons [15,16] and nucleons [21,22] extracted from $e^{-}e^{+}$ annihilation reactions. Finally, the developed approach can also be useful for analyzing fusion reactions at low energies.

### v3

@@ IV Conclusion
The effect of wave-like enhancement of the quantity $\sigma_{0}(E)v$ as $E\rightarrow 0$ was found by us in Mel in analysing fusion reactions (6) and (7). In the present work, we have investigated its manifestation in the experimental cross sections of $\Lambda\bar{\Lambda}$ production at $e^{-}e^{+}$ annihilation; however, in our opinion, it should be taken into account in any two-particle near-threshold reaction. It seems to us that a natural further development of the approach would be its generalization to describe the reaction of $\Lambda_{c}\bar{\Lambda}_{c}$ pair production at $e^{-}e^{+}$ annihilation Bes3 . However, it is necessary to include the Coulomb interaction between $\Lambda_{c}$ and $\bar{\Lambda}_{c}$ in the computational scheme, as well as to take into account the d-wave (3,4). The obtained formulas could be useful for analyzing the oscillatory behavior of the electromagnetic form factors of hyperons Heid2 ; Heid3 and nucleons Tomasi ; Yang extracted from $e^{-}e^{+}$ annihilation reactions. Finally, the developed approach can also be useful for analyzing fusion reactions at low energies.